OpenStudy (anonymous):

what is the solution set for |-2x+8|=12

OpenStudy (anonymous):

2x+8=12 2x = 4 x=2

OpenStudy (anonymous):

also... => -2x+8=+or - 12 =>x=(+12-8)/-2 =-2 or =>x=(-12-8)/-2=10 soluion set={-2,10}

OpenStudy (anonymous):

check for the solution sets by substituting them in the above equation

OpenStudy (anonymous):

Nice touch!